OpenStudy (anonymous):

Find the value of the constant c for which the line y=2x+c is a tangent to the curve y^2 =4x.

OpenStudy (anonymous):

see

OpenStudy (anonymous):

the idea is that you need to equate both equations as they intersect. Tangent gives equal roots so you both b^2-4ac=0

OpenStudy (anonymous):

It's the right step, but the answer says c= 1/2...?

OpenStudy (anonymous):

yes sorry i wrote -16c instead of -32c you will get now c=1/2

OpenStudy (anonymous):

why would it be -32c?

OpenStudy (anonymous):

Oh, I got it. Thanks :)

OpenStudy (anonymous):